a building casts a shadow 40 ft long when the sun's angle of elevation is 58 degrees. Find the height of the building.
Answer by mananth(16946)

a building casts a shadow 40 ft long when the sun's angle of elevation is 58 degrees. Find the height of the building.
...
let the height of building be x
tan(58 degrees) = x/40
1.60033453 = x/40
x = 1.60033453 * 40
x = 64.01 feet
...
m.ananth@hotmail.ca